Green Thumb Nursery has 37,000 shares outstanding at a market price of $63.25 per share. The earnings per share are $3.25. The firm has total assets of $330,000 and total liabilities of $191,000. Today, the firm announced a share repurchase for $85,000 of its stock. What is the earnings per share after the repurchase?
Answer #1

Total earnings=Number of shares*Earning per share =37000*3.25 =120250

Shares repurchased =amount of repurchase/current price =85000/63.25 =1343.87

Shares after repurchase =37000-1343.87 =35656.13

EPS after repurchase =120250/35656.13 =3.37